OpenStudy (anonymous):

Solve for x by using base 10 logs.

OpenStudy (anonymous):

+11^x+2 = 12^7x

OpenStudy (valpey):

Is this what you mean?: \[11^x+2=12^{7x}\]

OpenStudy (valpey):

Or\[11^x+2=12^7x\]

OpenStudy (valpey):

or \[11^{x+2}=12^{7x}\]

OpenStudy (anonymous):

yea but 11^(x+2) and the first one 12^(7X)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that last one

OpenStudy (valpey):

Okay (equation editor is for your use) \[11^{x+2}=12^{7x}\] \[\log_{10}{11^{x+2}}=\log_{10}12^{7x}\] \[\log_{10}{11^{x+2}}=\log_{10}12^{7x}\] \[(x+2)\log_{10}{11}=7x\log_{10}12\] \[(x+2)\log_{10}{11}=7x\log_{10}12\] \[2\log_{10}{11}=x(7\log_{10}12-\log_{10}{11})\] \[\frac{2\log_{10}{11}}{(7\log_{10}12-\log_{10}{11})}=x\] Does that look right?

OpenStudy (anonymous):

yep sounds right thnks :)

OpenStudy (valpey):

I don't know why it repeated itself there, but okay.